OpenStudy (anonymous):

what is the equation in standard form of a parallel line that passes through (1,-5)?

OpenStudy (mathmate):

Parallel to which line? Assuming the line has a given slope of "a" as in y=ax+b then the point-slope form of the line is \[(y-y_1)=a(x-x_1)\] where (x1,y1)=(1,-5)

OpenStudy (acxbox22):

\(\Huge\color{green}{\bigstar}\Huge\color{green}{\bigstar}\Huge\color{green}{\bigstar}\Large\color{orange}{WELCOME~TO~OPENSTUDY}\Huge\color{green}{\bigstar}\Huge\color{green}{\bigstar}\Huge\color{green}{\bigstar}\)

OpenStudy (acxbox22):

parallel means same slope different y intercept